OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (cj49):

let the three intergers be n,n+1,n+2 n+n+1+n+2=267 3n+3=267 3n=267-3 n=264/3 n=88 largest integer 88+2=90